# Dynamical Measurement of Supermassive Black Hole Masses: QPE Timing Method

### v2

@@ Abstract
pubyear: 2025pagerange: Dynamical Measurement of Supermassive Black Hole Masses: QPE Timing MethodA
@@ 4.1 GSN 069
In Jeffreys’ scale, disk precession in GSN 069 is substantially disfavored. 222 Miniutti et al. (2025) claimed that the super-orbital modulation in the GSN 069 data is not likely from EMRI apsidal precession, but is from disk precession. Probably the most reliable way to test this claim is a simultaneous measurement of the EMRI apsidal precession period $T_{\rm aps}$ and the disk precession period $\tau_{\rm p}$ from the data assuming EMRI+a precessing disk ($\mathcal{H}_{1}$). However, the measurement result $\tau_{\rm p}\gg T_{\rm aps}\approx 76$ day [Eq. (16)] supports the opposite: the disk precession is slow. NOTE that the $\tau_{\rm p}$ and $T_{\rm aps}$ constraints are measured from the GSN 069 data, instead of an assumption.
@@ Appendix A Posterior corner plots of GSN 069
In the Appendix, we include posterior corner plots of model parameters in hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ for analyzing QPE timing data of GSN 069.

### v3

@@ Abstract
pubyear: 2025pagerange: Dynamical Measurement of Supermassive Black Hole Masses: QPE Timing MethodC
@@ 4.1 GSN 069
In Jeffreys’ scale, disk precession in GSN 069 is substantially disfavored. 222 Miniutti et al. (2025) claimed that the super-orbital modulation in the GSN 069 data is not likely from EMRI apsidal precession, but is from disk precession. Probably the most reliable way to test this claim is a simultaneous measurement of the EMRI apsidal precession period $T_{\rm aps}$ and the disk precession period $\tau_{\rm p}$ from the data assuming EMRI+a precessing disk ($\mathcal{H}_{1}$). However, the measurement result $\tau_{\rm p}\gg T_{\rm aps}\approx 76$ day [Eq. (16)] supports the opposite: the disk precession is slow. NOTE that the $\tau_{\rm p}$ and $T_{\rm aps}$ constraints are measured from the GSN 069 data in the framework of EMRI+disk model, instead of an assumption. In Appendix B, we perform detailed O-C analyses for comparison and discuss possible sources of discrepancy in interpreting the super-orbital modulation in the GSN 069 data.
@@ Appendix A Posterior corner plots of GSN 069
In the section, we include posterior corner plots of model parameters in hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ for analyzing QPE timing data of GSN 069.
@@
Appendix B O-C analysis of GSN 069
For comparing with the Bayes analysis in the framework of EMRI+disk model, we apply a similar O-C analysis as in Miniutti et al. (2025) to GSN 069 data. In the top two panels of Fig. 9, we show the fitting result of a simple linear model, i.e., a constant period $T_{\rm trial}$,
${\rm O-C}=t_{\rm obs}^{(k)}-N_{\rm cyc}T_{\rm trial}\ ,$ (23)
where $N_{\rm cyc}$ is the number of cycles and $T_{\rm trial}=17.88$ hr (Miniutti et al., 2025). In the O-C analysis, there is some ambiguity in identifying $N_{\rm cyc}$ of each flare due to data gaps, while it is of no such difficulty in the Bayes analysis of the EMRI+disk model. Therefore, we use the numbers $N_{\rm cyc}$ from the best-fit EMRI+disk model in the O-C analysis. Note that the first flare in the last epoch is identified as the 213th flare, while it was identified as the 211th in Miniutti et al. (2025). From top panels of Fig. 9, the linear model of a constant period is insufficient to fit the data. In middle panels, we use a quadratic model
${\rm O-C_{quadratic}}=t^{(k)}_{\rm obs}-\left(T_{\rm obt}N_{\rm cyc}+\frac{1}{2}\dot{T}_{\rm obt}\langle T_{\rm obt}\rangle N_{\rm cyc}^{2}\right)\ ,$ (24)
where $\langle T_{\rm obt}\rangle$ is the averaged orbital period across the whole time span assuming constant $\dot{T}_{\rm obt}$. Fitting the quadratic model to either the odd and even flares, orbital period and its derivative are constrained as
[…]
and
[…]
at 2-$\sigma$ confidence level. The O-Cquadratic analysis does not yield clear in-phase or anti-phase pattern in the residuals of even and odd flares. In the bottom panels are the residuals of the best-fit QPE timing model with $\mathcal{H}_{0}$ for GSN 069, i.e. $t^{(k)}_{\rm obs}-t_{0}^{(k)}$, in which there is no clear signature of unmodeled super-orbital modulation either.
Another possible source of discrepancy is the different ways of defining the QPE timing: we define a starting time for each flare and interpret it as when the star crosses the disk, while Miniutti et al. (2025) use the flare peak time. We again apply the O-C analysis to flare peak times and the results are displayed in Fig. 10. The results are quit similar to what is show in Fig. 9 except lower uncertainties of the flare peak times.
To summarize, there is some intrinsic ambiguity for identifying number of cycles of each flare in O-C analysis. This may lead to different interpretations of O-C analysis results and is likely the source of discrepancy between this work and Miniutti et al. (2025).
Figure 9: O-C analysis versus EMRI+disk model fitting to GSN 069 data, where error bars represent 1-$\sigma$ uncertainties. The number of cycles of the first flare in the last epoch is labeled by the dashed vertical line.
Figure 10: Same to Fig. 9 except for that flare peak times instead of flare starting times are used.
Appendix C Best-fit EMRI trajectories disk motion
In this section, we present the best-fit EMRI trajectories with both $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ for the simulations in Section 4.3.